Find the following values. Round the answers to three decimal places. (Confidence Intervals for the Variance and Standard Deviation)

X^2(12,.025)

X^2(12,.975)

X^2(5,.005)

X^2(5,.995)

X^2(22,.1)

X^2(22,.9)
Answer #1

below are the corresponding value for chi square distribution:

X^2(12,.025) =4.404

X^2(12,.975) =23.337

X^2(5,.005) =0.412

X^2(5,.995) =16.750

X^2(22,.1) =14.041

X^2(22,.9) =30.813